# Cohort-Individual Cooperative Learning for Multimodal Cancer Survival Analysis

### v1

@@ Abstract
Recently, we have witnessed impressive achievements in cancer survival analysis by integrating multimodal data, e.g., pathology images and genomic profiles. However, the heterogeneity and high dimensionality of these modalities pose significant challenges for extracting discriminative representations while maintaining good generalization. In this paper, we propose a Cohort-individual Cooperative Learning (CCL) framework to advance cancer survival analysis by collaborating knowledge decomposition and cohort guidance. Specifically, first, we propose a Multimodal Knowledge Decomposition (MKD) module to explicitly decompose multimodal knowledge into four distinct components: redundancy, synergy and uniqueness of the two modalities. Such a comprehensive decomposition can enlighten the models to perceive easily overlooked yet important information, facilitating an effective multimodal fusion. Second, we propose a Cohort Guidance Modeling (CGM) to mitigate the risk of overfitting task-irrelevant information. It can promote a more comprehensive and robust understanding of the underlying multimodal data, while avoiding the pitfalls of overfitting and enhancing the generalization ability of the model. By cooperating the knowledge decomposition and cohort guidance methods, we develop a robust multimodal survival analysis model with enhanced discrimination and generalization abilities. Extensive experimental results on five cancer datasets demonstrate the effectiveness of our model in integrating multimodal data for survival analysis. The code will be publicly available soon.
@@ 1 Introduction
Survival analysis, one of the most important tasks of cancer prognosis, aims to assess the probability of an event (typically death in survival analysis) occurring for a particular patient and accurately rank the risks of cancer patients. It offers insights into disease progression, treatment effectiveness, and patient prognosis, ultimately leading to improved decision-making and patient care in research and clinical scenarios. However, the complex nature of cancer necessitates a comprehensive evaluation of diverse personalized data, posing a significant challenge for survival analysis models to effectively capture and incorporate this heterogeneity. Therefore, the development of an effective multimodal integration approach is essential yet challenging for constructing robust and accurate survival analysis models.
@@
Recent advances in Deep Learning (DL) [1, 2] have made survival analysis more efficient and accurate by leveraging patient’s clinical data [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], e.g., genomics and pathology images, reducing the workloads of clinicians considerably. Genomic profiles provide molecular information, enabling personalized medicine and understanding cancer genetics, while pathology images offer visual details that assist in diagnosis, grading, and assessing tumor heterogeneity. Together, these high-dimensional data enhance the understanding of tumours, thereby facilitating advancements in patient care and cancer research. Recently, DL-based multimodal methods [18, 19, 20, 21, 22, 23, 24] integrated these two modalities to enable the acquisition of complementary information from diverse perspectives, promising more accurate survival analysis. For example, several works [25, 26, 27, 28, 29, 30] focused on enhancing modality representations by leveraging cross-modal interactions. However, the challenges posed by the heterogeneity gap are still not well addressed, thus undermining the efficacy of multimodal integration. Furthermore, the high dimensionality increases the risk of overfitting task-irrelevant information, resulting in performance degradation on unseen samples.
@@
By tackling the above issues, our goal is to construct a more general and effective multimodal survival analysis model by incorporating a comprehensive knowledge decomposition and a more general patient cohort guidance. First, the importance of different knowledge components in the integration process is variable. For example, common knowledge shared by multiple modalities typically is redundant in multimodal integration, thus interfering models to learn other discriminative information. Furthermore, synergy is the new knowledge generated from multimodal interactions and may be overlooked without explicit modeling. In our framework, we present a comprehensive decomposition of multimodal knowledge into four distinct components: redundancy, synergy, and uniqueness of the two modalities. This decomposition will enable a deeper understanding of underlying factors impacting survival outcomes. Second, extracting discriminative features from high-dimensional data while ensuring good generalization is a tough challenge. For example, a wealth of task-irrelevant information can produce spurious correlations between modalities, demanding elaborated solutions to learn multimodal interactions with enhanced generalization ability. Therefore, we seek to develop a more general patient cohort guidance mechanism, allowing for a broader range of patient characteristics to be considered during model training. Such guidance can prevent the models from overemphasizing task-irrelevant information and thus enhance the model’s generalization ability. By combining these advancements, we can construct an effective and robust multimodal survival analysis model with enhanced accuracy and applicability.
@@
In this paper, we propose a Cohort-individual Cooperative Learning (CCL) framework to integrate genomics and pathology images for cancer survival analysis. Specifically, first, we propose a Multimodal Knowledge Decomposition (MKD) module to decompose multimodal knowledge into four distinct components: redundancy, synergy, and uniqueness of the two modalities. Such a comprehensive decomposition serves as an illuminating framework, enabling models to discern often disregarded yet crucial information. It paves the way for effective multimodal fusion, enhancing the integration of diverse data modalities and their complementary information. Second, we propose a Cohort Guidance Modeling (CGM) to unleash the potential of distinct knowledge components and to enhance the model’s generalization ability. Our cohort guidance assists feature learning to at both knowledge and patient levels, capturing the essence of multifaceted data at various levels of granularity. By cooperating knowledge decomposition and cohort guidance, we enhance our model’s discrimination and generalization abilities by effectively fusing diverse modalities while mitigating overfitting risks. Experiment results on five datasets in The Cancer Genome Atlas (TCGA) program prove that our framework achieves state-of-the-art performance in survival analysis.
@@ 2.1 Unimodal Survival Analysis
Survival analysis can be expressed as an estimation of the hazard function, which models the patient’s probability of death at a certain time, conditioned on personal clinical records. In an early stage, Cox’s proportional hazards regression model [31] conceptualises the hazard function as a multiplication of two components: 1) the underlying baseline hazard function, describing how the risk of event per time unit changes over time at baseline levels of covariates; and 2) the hazard ratio, measuring the impact of covariates. Since the baseline hazard function for a given cancer is constant, the hazard ratio determines whether a patient is at high risk. Based on the Cox’s regression model, subsequent survival analysis methods [3, 4, 5, 7, 6, 8] aim to predict the personalized hazard ratio using quantitative data extracted from short-term clinical indicators or long-term follow-up reports. For example, Kappen et al. [6] summarized seventeen pretreatment characteristics, such as residual tumor size, age, and thrombocytes, to predict the treatment outcome using a neural network. Ohno-Machado et al. [8] predicted the survival of AIDS patients conditioned on demographics, laboratory markers, and clinical findings, such as age, hemoglobin, and albumin.
Genomic profiles provide molecular information of tumors, which is important for cancer prognosis prediction. For example, certain genetic mutations or variations can affect tumor growth, metastasis, and response to chemotherapy or targeted therapies. Extensive studies [9, 10, 12, 11] have been conducted to predict cancer prognosis and treatment response by leveraging genomics data. Recently, Yousefi et al. [12] combined deep learning and Bayesian optimization methods to tackle high-dimensional cancer outcomes prediction tasks. Qiu et al. [11] built good predictive models with limited high-dimensional samples by using a meta-learning survival analysis framework.
@@
Pathological images provide morphological features of tumors, which can provide valuable information about the aggressiveness of the tumor, its response to treatment, and the likelihood of disease recurrence. Some recent works [13, 14, 15, 16] constructed effective survival analysis models based on giga-pixel Whole Slide Images (WSIs). For example, Yao et al. [15] introduced the siamese MI-FCN and attention-based MIL pooling to efficiently learn imaging features from the WSI and then aggregate WSI-level information to patient-level. Zhu et al. [16] adaptively sampled and grouped hundreds of patches from each WSI into different several clusters. Then they employed an aggregation model to make patient-level predictions based on cluster-level survival prediction results.
@@ 2.2 Multimodal Survival Analysis
Integrating genomics and pathology images can enhance the predictive power of survival analysis models [18, 19, 20, 32, 25, 22, 21, 24, 23] and thus attract increasing attention recently. By utilizing the cellular graph embedded in the tissue, Nakhli et al. [22] utilized a unified representation for each patient, leveraging the hierarchical organization of the tissue. Chen et al. [21] captured the interactions between features across multiple modalities by combining the unimodal feature representations using the Kronecker product. Additionally, they incorporated a gating-based attention mechanism to control the expressive power of each representation. Furthermore, Chen et al. [25] proposed an interpretable, dense co-attention mapping between WSIs and genomic features formulated in the embedding space. Xu et al. [28] introduced the optimal transport theory to match WSI patches and gene embeddings for selecting informative patches to represent gigapixel WSIs and build an interpretable co-attention module to effectively fuse multimodal data. Moreover, Zhou et al. [29] found that the generated cross-modal representation can enhance and recalibrate intra-modal representation, and thus significantly improve its discrimination for survival analysis.
@@ 3.1 Unimodal Feature Extraction
Genomics. Genomic profiles can identify specific genetic alterations or biomarkers associated with cancer prognosis. For example, certain genetic mutations, gene expression patterns, or alterations in DNA copy number can serve as prognostic markers, helping to predict the likelihood of survival conditions of patients. In our framework, we partition RNA sequencing (RNA-seq), Copy Number Variation (CNV), Simple Nucleotide Variation (SNV) sequences into six sub-sequences as previous methods [29, 28]. Each sub-sequence is transformed into a feature by a Self-normalizing Neural Network (SNN) [17], which is a more powerful alternative to conventional Multi-Layer Perceptron (MLP). By integrating all sub-sequence features, we obtain the genomic representation $F_{g}\in\mathbb{R}^{1\times 256}$.
@@
Pathology. Whole Slide Images (WSIs), i.e., pathology images, describe the information about the tumor immune microenvironment and provide valuable information for cancer prognosis prediction. Considering the huge resolution of WSI, which exceeds the capacity of Convolutional Neural Networks (CNNs), we adopt a strategy of splitting tissue regions within each WSI into non-overlapping patches at 20x magnification (256 x 256 resolution for each patch). Following previous works [28, 29, 21], we utilize an ImageNet pre-trained ResNet-50 [2] model to extract a 1024-dimensional embedding for each patch, while all patch embeddings of the same WSI are collected as an embedding set. It is worth noting that such an embedding set is still high-dimensional as it typically contains tens of thousands of patches for each WSI. To further reduce information redundancy, we employ the K-means algorithm [33] to cluster all patch embeddings into $k$ groups and leverage the cluster centers as pathology features. However, due to the stochastic nature of K-means, cluster centers of different samples may be misaligned, which means that two cluster centers from the same ordinal position of two samples may exhibit completely different phenotypes. Consequently, deep models prefer to learn phenotype-independent knowledge instead of specialized knowledge tailored to each phenotype, overlooking crucial information in important patches. In our framework, we address this misalignment issue by cluster center alignment (CCA), an optimal matching between an anchor and cluster centers, as shown in Fig. 3. This involves assigning each cluster center to a feature in the anchor with maximized similarities of matched pairs. To solve this, the Hungarian algorithm [34] is employed to calculate the permutation matrix that maps the current centers to their matched ordinal positions. After that, we obtained the aligned centers by multiplying cluster centers with the permutation matrix. Aligned centers are utilized to update the anchor with a ratio of $\tau$ during training and generated the pathology representation $F_{p}\in\mathbb{R}^{1\times 256}$, ensuring accurate alignment and capturing crucial information.
@@
Figure 3: Cluster center alignment for pathology features. Aligned centers at the same positions typically exhibit similar phenotypes across patients, enabling the network to extract specific information pertaining to these phenotypes.
@@ 3.2 Cohort-individual Cooperative Learning
Multimodal Knowledge Decomposition. In our framework, we decompose the knowledge within genomic profiles and pathology images into four distinct components, including redundancy, synergy and uniqueness of two modalities. Concretely, the large heterogeneity gap between pathology and genomics indicates a wealth of modality-specific knowledge. Meanwhile, there still exist considerable common knowledge between them. For example, particular genetic mutations or alterations can be observed at both the molecular level and the histopathological level simultaneously. Furthermore, another essential component, synergy, can provide new insights on tumors to enrich the knowledge in multimodal data, especially for heterogeneous modalities. Emerging from multimodal interactions, it surpasses common knowledge by forming a unique realm of insights unattainable through any individual modality alone. As an example, the diagnosis of glioma typically necessitates genetic markers (IDH mutation and 1p/19q co-deletion) to categorize gliomas into subtypes and histopathological images to determine the presence of microvascular proliferation and necrosis simultaneously. Overall, a comprehensive and explicit decomposition enhances integration, representation, synergy capture, and flexibility. It enables more effective modeling and utilization of multimodal data, leading to improved performance, insights, and decision-making in cancer prognosis.
To achieve a comprehensive decomposition of knowledge, we develop four encoders within our framework. These encoders are specifically employed to capture and model different components—redundancy, synergy, and uniqueness—present within the genomic profiles and pathology images. Concretely, we employ MLP layers as modality encoders $\Phi_{p}$ and $\Phi_{g}$ to extract specific knowledge $P$ and $G$, focusing on aspects distinct from common knowledge shared by both modalities. The formula can be written as:
@@
$\centering C=\Phi_{c}(F_{p},F_{g})\@add@centering$ $=fc(A^{T})*F_{p}+fc(A)*F_{g},$ (2)
$A$ $=fc(F_{p})^{T}fc(F_{g}),$ (3)
where $A$ is the co-attention matrix and $fc$ indicates a fully-connected layer. Synergy $S$ is computed the same as common knowledge $C$, but with different parameters. Using the above encoders, we decompose modality representations $F_{g}$ and $F_{p}$ into genomic-specific $G$, pathology-specific $P$, common $C$, and synergistic $S$ components, respectively.
Cohort Guidance Modeling. In the above section, we employ different encoders to capture different knowledge components from multimodal data. Beside that, an effective loss function is also significant in enhancing the discrimination and generalization abilities of our model. Due to the high dimensionality of input modalities, multimodal interactions learned from each patient, as done in existing methods [25, 28, 29], may overfit task-irrelevant information, leading to reduced generalization and discrimination abilities. Therefore, we propose to utilize the correlations between patient cohorts to acquire more general knowledge components with cohort consistency, capture the heterogeneity of multimodal data, and gain a better understanding of cancer.
Figure 4: The structure of common and synergistic encoders. We generate attention vectors based on the co-attention matrix to integrate multimodal features. The learnable parameters for common and synergistic knowledge are different, enabling the same structure used to extract distinct knowledge components.
In our framework, we harness the cohort guidance to learn knowledge components from both knowledge and patient levels, as shown in Fig. 5. At the knowledge level, the most prominent difference between the decomposed knowledge components is their correlation to the original modality representations. Specifically, synergy is unattainable through any individual modality alone, so it is beyond the knowledge of both modalities. Redundancy is the intersection of knowledge as it is shared by both modalities. In addition, uniqueness is located in the source modality and is distinct from another modality. Therefore, we employ a set of similarity constrains between them as:
@@
where $cos$ computes the cosine similarity between two inputs and ||$\cdot$|| ⋅ | is the absolute operation used to constraints two inputs orthogonal to each other. We freeze the gradients of $F_{p}$ and $F_{g}$ from $l_{k}$ to prevent potential model collapse that the extracted pathology and genomic features follow similar distributions. At the patient level, task-relevant information can be obtained by distinguishing patients conditioned on their risk scores, which are valuable for extracting more discriminative representations. Therefore, we split all patients into $r$ equal groups conditioned on their ground truth survival time, while patients in the same group are considered as similar. For uncensored patient, its features are closer to the features extracted from similar patients, while distinctive to patients in other groups. For censored patients, it may be closer to the features of patients at similar or lower risk and distinctly different from those at higher risk. Such an assumption in line with the definition of contrastive learning, which has been proved to have strong feature learning capability in both supervised and unsupervised settings. Therefore, we develop a cohort contrastive learning to enhance the discrimination and generalization abilities of our model. In the training phase, we build a cohort bank to store features of historical patients in each group. Our cohort bank is actually $r$ queues with length $b$ and follows the first-in-first-out principle for dynamically feature updating. For the feature of uncensored patient, we reduce its distances to similar patients in the cohort bank, while enlarging its distances to other patients. As an example, the formula for synergy $S$ is:
@@
where $\mathcal{S}_{+}$ is the patient group of similar risk storing in the cohort bank, while $\mathcal{S}_{-}$ is the collection of other patients. $d$ is the similarity function. Noted that $S$ can be replaced by other knowledge components $G$, $P$, or $C$ similarly. For censored patient, we can extend $\mathcal{S}_{+}$ as the collection of patients with similar or lower risk.

### v2

@@ Abstract
Recently, we have witnessed impressive achievements in cancer survival analysis by integrating multimodal data, e.g., pathology images and genomic profiles. However, the heterogeneity and high dimensionality of these modalities pose significant challenges in extracting discriminative representations while maintaining good generalization. In this paper, we propose a Cohort-individual Cooperative Learning (CCL) framework to advance cancer survival analysis by collaborating knowledge decomposition and cohort guidance. Specifically, first, we propose a Multimodal Knowledge Decomposition (MKD) module to explicitly decompose multimodal knowledge into four distinct components: redundancy, synergy, and uniqueness of the two modalities. Such a comprehensive decomposition can enlighten the models to perceive easily overlooked yet important information, facilitating an effective multimodal fusion. Second, we propose a Cohort Guidance Modeling (CGM) to mitigate the risk of overfitting task-irrelevant information. It can promote a more comprehensive and robust understanding of the underlying multimodal data while avoiding the pitfalls of overfitting and enhancing the generalization ability of the model. By cooperating with the knowledge decomposition and cohort guidance methods, we develop a robust multimodal survival analysis model with enhanced discrimination and generalization abilities. Extensive experimental results on five cancer datasets demonstrate the effectiveness of our model in integrating multimodal data for survival analysis. Our code is available at https://github.com/moothes/CCL-survival.
@@ 1 Introduction
Survival analysis, one of the most important tasks of cancer prognosis, aims to assess the probability of an event (typically death in survival analysis) occurring for a particular patient and accurately rank the risks of cancer patients. It offers insights into disease progression, treatment effectiveness, and patient prognosis, ultimately leading to improved decision-making and patient care in research and clinical scenarios [1]. However, the complex nature of cancer necessitates a comprehensive evaluation of diverse personalized data, posing a significant challenge for survival analysis models to capture and incorporate this heterogeneity effectively. Therefore, developing an effective multimodal integration approach is essential yet challenging for constructing robust and accurate survival analysis models.
@@
Recent advances in Deep Learning (DL) [2, 3, 4] have made survival analysis more efficient and accurate by leveraging patients’ clinical data [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], e.g., genomics and pathology images, reducing clinicians’ workloads considerably. Genomic profiles provide molecular information, enabling personalized medicine and understanding of cancer genetics, while pathology images offer visual details that assist in diagnosis, grading, and assessing tumor heterogeneity. Together, these high-dimensional data enhance the understanding of tumors, thereby facilitating advancements in patient care and cancer research. Recently, DL-based multimodal methods [21, 22, 23, 24, 25, 26, 27, 28] integrated these two modalities to enable the acquisition of complementary information from diverse perspectives, promising more accurate survival analysis. For example, several works [29, 30, 31, 32, 33, 34] focused on enhancing modality representations by leveraging cross-modal interactions. However, the challenges posed by the heterogeneity gap are still not well addressed, thus undermining the efficacy of multimodal integration. Furthermore, the high dimensionality increases the risk of overfitting task-irrelevant information, resulting in performance degradation on unseen samples.
@@
By tackling the above issues, our goal is to construct a more general and effective multimodal survival analysis model by incorporating a comprehensive knowledge decomposition and a more general patient cohort guidance. First, the importance of different knowledge components in the integration process is variable. For example, common knowledge shared by multiple modalities typically is redundant in multimodal integration, thus interfering with models to learn other discriminative information. Furthermore, synergy is the new knowledge generated from multimodal interactions and may be overlooked without explicit modeling. In our framework, we present a comprehensive decomposition of multimodal knowledge into four distinct components: redundancy, synergy, and uniqueness of the two modalities. This decomposition will enable a deeper understanding of underlying factors impacting survival outcomes. Second, extracting discriminative features from high-dimensional data while ensuring good generalization is a tough challenge. For example, a wealth of task-irrelevant information can produce spurious correlations between modalities, demanding elaborated solutions to learn multimodal interactions with enhanced generalization ability. Therefore, we seek to develop a more general patient cohort guidance mechanism, allowing for a broader range of patient characteristics to be considered during model training. Such guidance can prevent the models from overemphasizing task-irrelevant information and thus enhance the model’s generalization ability. By combining these advancements, we can construct an effective and robust multimodal survival analysis model with enhanced accuracy and applicability.
@@
In this paper, we propose a Cohort-individual Cooperative Learning (CCL) framework to integrate genomics and pathology images for cancer survival analysis. Specifically, first, we propose a Multimodal Knowledge Decomposition (MKD) module to decompose multimodal knowledge into four distinct components: redundancy, synergy, and uniqueness of the two modalities. Such a comprehensive decomposition serves as an illuminating framework, enabling models to discern often disregarded yet crucial information. It paves the way for effective multimodal fusion, enhancing the integration of diverse data modalities and their complementary information. Second, we propose a Cohort Guidance Modeling (CGM) to unleash the potential of distinct knowledge components and to enhance the model’s generalization ability. Our cohort guidance assists feature learning to at both knowledge and patient levels, capturing the essence of multifaceted data at various levels of granularity. By cooperating with knowledge decomposition and cohort guidance, we enhance our model’s discrimination and generalization abilities by effectively fusing diverse modalities while mitigating overfitting risks. Experiment results on five datasets in The Cancer Genome Atlas (TCGA) program prove that our framework achieves state-of-the-art performance in survival analysis.
@@ 2.1 Unimodal Survival Analysis
Survival analysis can be expressed as an estimation of the hazard function, which models the patient’s probability of death at a certain time, conditioned on personal clinical records. In an early stage, Cox’s proportional hazards regression model [35] conceptualizes the hazard function as a multiplication of two components: 1) the underlying baseline hazard function, describing how the risk of event per time unit changes over time at baseline levels of covariates; and 2) the hazard ratio, measuring the impact of covariates. Since the baseline hazard function for a given cancer is constant, the hazard ratio determines whether a patient is at high risk. Based on Cox’s regression model, subsequent survival analysis methods [5, 6, 7, 9, 8, 10] aim to predict the personalized hazard ratio using quantitative data extracted from short-term clinical indicators or long-term follow-up reports. For example, Kappen et al. [8] summarized seventeen pretreatment characteristics, such as residual tumor size, age, and thrombocytes, to predict the treatment outcome using a neural network. Ohno-Machado et al. [10] predicted the survival of AIDS patients based on demographics, laboratory markers, and clinical findings, such as age, hemoglobin, and albumin.
Genomic profiles provide molecular information on tumors, which is important for cancer prognosis prediction. For example, certain genetic mutations or variations can affect tumor growth, metastasis, and response to chemotherapy or targeted therapies. Extensive studies [11, 12, 14, 13] have been conducted to predict cancer prognosis and treatment response by leveraging genomics data. Recently, Yousefi et al. [14] combined deep learning and Bayesian optimization methods to tackle high-dimensional cancer outcomes prediction tasks. Qiu et al. [13] built good predictive models with limited high-dimensional samples by using a meta-learning survival analysis framework.
@@
Pathological images provide morphological features of tumors, which can provide valuable information about the aggressiveness of the tumor, its response to treatment, and the likelihood of disease recurrence. Some recent works [15, 16, 17, 18] constructed effective survival analysis models based on giga-pixel Whole Slide Images (WSIs). For example, Yao et al. [17] introduced the siamese MI-FCN and attention-based MIL pooling to efficiently learn imaging features from the WSI and then aggregate WSI-level information to the patient level. Zhu et al. [18] adaptively sampled and grouped hundreds of patches from each WSI into several clusters. Then they employed an aggregation model to make patient-level predictions based on cluster-level survival prediction results.
@@ 2.2 Multimodal Survival Analysis
Integrating genomics and pathology images can enhance the predictive power of survival analysis models [21, 22, 23, 36, 26, 24, 28, 27, 37] and thus attract increasing attention recently. By utilizing the cellular graph embedded in the tissue, Nakhli et al. [26] extracted a unified representation for each patient, leveraging the hierarchical organization of the tissue. Chen et al. [24] captured the interactions between features across multiple modalities by combining the unimodal feature representations using the Kronecker product. Additionally, they incorporated a gating-based attention mechanism to control the expressive power of each representation. Furthermore, Chen et al. [29] proposed an interpretable, dense co-attention mapping between WSIs and genomic features formulated in the embedding space. Xu et al. [32] introduced the optimal transport theory to match WSI patches and gene embeddings for selecting informative patches to represent gigapixel WSIs and build an interpretable co-attention module to effectively fuse multimodal data. Moreover, Zhou et al. [33] found that the generated cross-modal representation can enhance and recalibrate intra-modal representation, and thus significantly improve its discrimination for survival analysis.
@@ 3.1 Unimodal Feature Extraction
Genomics. Genomic profiles can identify specific genetic alterations or biomarkers associated with cancer prognosis. For example, certain genetic mutations, gene expression patterns, or alterations in DNA copy number can serve as prognostic markers, helping to predict the likelihood of survival conditions of patients. In our framework, we partition RNA sequencing (RNA-seq), Copy Number Variation (CNV), and Simple Nucleotide Variation (SNV) sequences into six sub-sequences as previous methods [33, 32]. Each sub-sequence $g_{i}$ is transformed into a feature by two cascaded 256-dimension Self-normalizing Neural Network (SNN) [19] layers. Finally, we obtain the genomic representation $F_{g}=fc_{1}(SNN(SNN(g_{i})))$ by using a fully-connected layer $fc_{1}$ to aggregate the representations of multiple sequences into a single genomic representation $F_{g}\in\mathbb{R}^{1\times 256}$.
@@
Pathology. Whole Slide Images (WSIs), i.e., pathology images, describe the information about the tumor immune microenvironment and provide valuable information for cancer prognosis prediction. Considering the huge resolution of WSI, which exceeds the capacity of Convolutional Neural Networks (CNNs), we adopt a strategy of splitting tissue regions within each WSI into non-overlapping patches at 20x magnification (256 x 256 resolution for each patch). Following previous works [32, 33, 24], we utilize an ImageNet pre-trained ResNet-50 [3] model to extract a 1024-dimensional embedding for each patch, while all patch embeddings of the same WSI are collected as an embedding set. It is worth noting that such an embedding set is still high-dimensional as it typically contains tens of thousands of patches for each WSI. To further reduce information redundancy, we employ the K-means algorithm [38] to cluster all patch embeddings into $k$ groups and leverage the cluster centers as pathology features. However, due to the stochastic nature of K-means, cluster centers of different samples may be misaligned, which means that two cluster centers from the same ordinal position of two samples may exhibit completely different phenotypes. Consequently, deep models prefer to learn phenotype-independent knowledge instead of specialized knowledge tailored to each phenotype, overlooking crucial information in important patches. In our framework, we address this misalignment issue by cluster center alignment (CCA), an optimal matching between an anchor and cluster centers, as shown in Fig. 3. This involves assigning each cluster center to a feature in the anchor with maximized similarities of matched pairs. To solve this, the Hungarian algorithm [39] is employed to calculate the permutation matrix that maps the current centers to their matched ordinal positions. After that, we obtain the aligned centers by multiplying cluster centers with the permutation matrix. The aligned centers are utilized to update the anchor with a ratio of $\tau$ during training. Finally, a 256-dimension SNN layer and a fully-connected layer $fc_{1}$ are employed to aggregate the cluster centers into a single pathology representation $F_{p}\in\mathbb{R}^{1\times 256}$. Overall, the pathology representation is generated by $F_{p}=fc_{1}(SNN(Matching(Kmeans(ResNet(p_{i})),B)))$, where $p_{i}$ is the pathology patch and $B$ is the anchor.
@@
Figure 3: Cluster center alignment for pathology features. Aligned centers at the same positions typically exhibit similar phenotypes across patients, enabling the network to extract specific information about these phenotypes.
@@ 3.2 Cohort-individual Cooperative Learning
Multimodal Knowledge Decomposition. In our framework, we decompose the knowledge within genomic profiles and pathology images into four distinct components, including redundancy, synergy, and uniqueness of two modalities. Concretely, the large heterogeneity gap between pathology and genomics indicates a wealth of modality-specific knowledge. Meanwhile, there still exists considerable common knowledge between them. For example, particular genetic mutations or alterations can be observed at both the molecular level and the histopathological level simultaneously. Furthermore, another essential component, synergy, can provide new insights inon tumors to enrich the knowledge of multimodal data, especially for heterogeneous modalities. Emerging from multimodal interactions, it surpasses common knowledge by forming a unique realm of insights unattainable through any individual modality alone. As an example, the diagnosis of glioma typically necessitates genetic markers (IDH mutation and 1p/19q co-deletion) to categorize gliomas into subtypes and histopathological images to determine the presence of microvascular proliferation and necrosis simultaneously. Overall, a comprehensive and explicit decomposition enhances integration, representation, synergy capture, and flexibility. It enables more effective modeling and utilization of multimodal data, leading to improved performance, insights, and decision-making in cancer prognosis.
To achieve a comprehensive knowledge decomposition, we develop four encoders within our framework. These encoders are specifically employed to capture and model different components—redundancy, synergy, and uniqueness [42, 43, 44, 45]—present within the genomic profiles and pathology images. Concretely, we employ MLP layers as modality encoders $\Phi_{p}$ and $\Phi_{g}$ to extract specific knowledge $P$ and $G$, focusing on aspects distinct from common knowledge shared by both modalities. The formula can be written as:
@@
$\centering C=\Phi_{c}(F_{p},F_{g})\@add@centering$ $=fc_{1}(A^{T})*F_{p}+fc_{1}(A)*F_{g},$ (2)
$A$ $=fc(F_{p})^{T}\cdot fc(F_{g}),$ (3)
where $A$ is the co-attention matrix, while $fc$ is a 256-dimension fully-connected layer to further fuse input features. $fc_{1}$ is a fully-connected layer that transforms $256\times 256$ co-attention matrix into a $1\times 256$ attention vector for pathology and genomics representations, respectively. The $\cdot$ means matrix multiplication, while * indicates element-wise multiplication. Synergy $S$ is computed the same as common knowledge $C$, but with different parameters. Using the above encoders, we decompose modality representations $F_{g}$ and $F_{p}$ into genomic-specific $G$, pathology-specific $P$, common $C$, and synergistic $S$ components, respectively.
Figure 4: The structure of common and synergistic encoders. We generate attention vectors based on the co-attention matrix to integrate multimodal features. The learnable parameters for common and synergistic knowledge are different, enabling the same structure used to extract distinct knowledge components. The * and $+$ indicate element-wise multiplication and addition, respectively.
Cohort Guidance Modeling. In the above section, we utilize various encoders to capture diverse knowledge components from multimodal data. Nonetheless, relying solely on different modules does not guarantee the effective extraction of the intended knowledge components. To overcome this challenge, the incorporation of additional supervision signals becomes necessary. These signals serve to guide multiple encoders in acquiring distinct knowledge components, thereby facilitating the process of multimodal fusion. Moreover, due to the high dimensionality of input modalities, multimodal interactions learned from each patient, as done in existing methods [29, 32, 33], may overfit task-irrelevant information, leading to reduced generalization and discrimination abilities. Based on the above analysis, we propose to utilize the correlations between patient cohorts to acquire more general knowledge components with cohort consistency, capture the heterogeneity of multimodal data, and gain a better understanding of cancer. In our framework, we harness the cohort guidance to learn knowledge components from both knowledge and patient levels, as shown in Fig. 5.
At the knowledge level, the most prominent difference between the decomposed knowledge components is their correlation to the original modality representations. Specifically, synergy is unattainable through any individual modality alone, so it is beyond the knowledge of both modalities. Redundancy is the intersection of knowledge as it is shared by both modalities. In addition, uniqueness is located in the source modality and is distinct from another modality. Therefore, we employ a set of similarity constraints between them:
@@
where $cos$ computes the cosine similarity between two inputs and ||$\cdot$|| ⋅ | is the absolute operation used to constrain two inputs orthogonal to each other. By minimizing $l_{k}$, multiple encoders are guided to learn distinct knowledge components. In this context, the genomic-specific component $G$ captures knowledge specifically derived from genomics data, distinct from the knowledge obtained from pathology data. To achieve this, we maximize the similarity between the genomic-specific component $G$ and the genomics representation $F_{g}$ by using the term $-cos(G,F_{g})$. Simultaneously, we enforce orthogonality between $G$ and the pathology representation $F_{p}$ by using the term $|cos(G,F_{p})|$. Similarly, the pathology-specific component $P$ follows a similar guidance strategy as $G$ to capture knowledge specific to pathology data. Furthermore, there exists common knowledge shared by both modalities. To model this shared knowledge, we maximize its similarities to both the genomics representation $F_{g}$ and the pathology representation $F_{p}$ by using the term $-cos(C,F_{p})-cos(C,F_{g})$. In addition, the synergistic component $S$ is supervised to minimize its similarities to both the genomics and pathology representations. Since it represents knowledge beyond the scope of both modalities, we use the term $|cos(S,F_{p})|+|cos(S,F_{g})|$ to achieve this objective. We freeze the gradients of $F_{p}$ and $F_{g}$ from $l_{k}$ to prevent the potential model collapse that the extracted features follow similar distributions.
At the patient level, task-relevant information can be obtained by distinguishing patients conditioned on their risk scores, which are valuable for extracting more discriminative representations. Therefore, we split all patients into $r$ equal groups conditioned on their ground truth survival time, while patients in the same group are considered as similar. For uncensored patients, its features are closer to the features extracted from similar patients, while distinctive to patients in other groups. For censored patients, it may be closer to the features of patients at similar or lower risk and distinctly different from those at higher risk. Such an assumption is in line with the definition of contrastive learning, which has been proven to have strong feature learning capability in both supervised and unsupervised settings. Therefore, we develop a cohort contrastive learning to enhance the discrimination and generalization abilities of our model. In the training phase, we build a cohort bank to store features of historical patients in each group. The cohort bank is implemented as a queue, following the “First In, First Out” (FIFO) principle. Specifically, given the bank length $b$ and current iteration $t$, the cohort bank stores the decomposed components from iteration ($t-b$) to ($t-1$). After the current iteration, the cohort bank will be updated to store the decomposed components from iteration ($t-b+1$) to ($t$). It is noteworthy that when $t<b+1$, no representations in the cohort bank will be discarded when updating.
For the feature of uncensored patients, we reduce its distances to similar patients in the cohort bank, while enlarging its distances to other patients. As an example, the formula for synergy $S$ is:
@@
where $\mathcal{S}_{+}$ is the patient group of similar risk storing in the cohort bank, while $\mathcal{S}_{-}$ is the collection of other patients. $d$ is the similarity function. Note that $S$ can be replaced by other knowledge components $G$, $P$, or $C$ similarly. For censored patients, we can extend $\mathcal{S}_{+}$ as the collection of patients with similar or lower risk.